OpenStudy (anonymous):

Find the area inside of r = 3 + 3sin(theta)

zepdrix (zepdrix):

Umm ok I'm trying to remember how to do these... Let's see.. So the formula for area in polar is:\[\large \int\limits_a^b \frac{1}{2}r^2\;d \theta\]Right? Soooooo, we're given our function r. Were you able to find the limits of integration? We could graph it, that might help.

OpenStudy (anonymous):

It should be 27pi/2 and I'm trying to solve it using a double integral. So it will be: \[\int\limits_{a}^{b}\int\limits_{f(\theta)}^{g(\theta)}rdrd \theta\] because the volume element, dA, is r dr dtheta, in polar coordinates

OpenStudy (anonymous):

I'm just having trouble finding the limits

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

EDIT: earlier I said 'volume element' I meant area element

zepdrix (zepdrix):

I think what we want to do is..... We want the area to sweep around from that origin point. So we would START from an angle of \(\large \theta=-\pi/2\) and sweep around and END at \(\large \theta=3\pi/2\)

zepdrix (zepdrix):

You can see that the distance we cover is 2pi. But I think if we integrate from 0 to 2pi, we might run into trouble. I'm not sure though :o I'd have to try it both ways to see...

OpenStudy (anonymous):

then the limits on r would be 0 to 3 + 3sin(theta) ?

zepdrix (zepdrix):

Our integral is being taken with respect to theta, so the limits of integration are with respect to theta. \[\huge \int\limits\limits_{\theta=a}^b \frac{1}{2}r^2\;d \theta\] Where \(\large a=-\pi/2\) and \(\large b=3\pi/2\) in this case. Maybe I misunderstood what you were saying though. Limits on r? Hmm. We don't have anything to subtract from r. But yes as you said we're going from 0 to our function r. If you wanted to, you could write your function as [3+3sin(theta)-0]. But blah.

zepdrix (zepdrix):

\[\huge \int\limits\limits\limits_{\theta=-\pi/2}^{3\pi/2} \frac{1}{2}(3+3\sin \theta)^2\;d \theta\]

OpenStudy (anonymous):

That results in 3pi. Also, I know it can be done using double integrals.

zepdrix (zepdrix):

3pi? ok maybe i made a mistake somewhere, gimme a sec :)

OpenStudy (anonymous):

But I'm trying to use this method: http://tutorial.math.lamar.edu/Classes/CalcIII/DIPolarCoords.aspx

OpenStudy (anonymous):

I understand how it works but I don't know how to setup the limits in this example.

OpenStudy (anonymous):

Ex 2 from that site is very similar. But when integrating from r=0 to r=3+3sin(theta), I don't know what the theta limits should be.

zepdrix (zepdrix):

Where are you getting 3pi from? That sounds way off...

OpenStudy (anonymous):

I'm doing all the integration in matlab. But I've solved it now.

OpenStudy (anonymous):

I've been literally mistyping the 3 + 3sin(theta) limit EVERYTIME as 3sin(theta) XD!!!

zepdrix (zepdrix):

Oh lololol

zepdrix (zepdrix):

If you were confused about the way they got the limits of integration, we can do the same thing as in the example ~ Find points of intersection. When does r=0 `intersect` r=3+3sin(theta) ? If you've got this all figured out already then no problem. Heh.

OpenStudy (anonymous):

So the solution is: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{3 + 3\sin(\theta)}r dr d \theta\] Thanks for your time though! You've kept me looking at this problem long enough to notice my mistake, and I appreciate this!

OpenStudy (anonymous):

I think I need some sleep... Thanks again!

zepdrix (zepdrix):

Mmmm that 0 and 2pi don't look correct :[ I think those would give you a problem if you use them for your limits on theta...

zepdrix (zepdrix):

But whatever, go to sleep if you're tired! lol